An investor will invest a total of $15000 in two accounts, one paying 4% annual simple interest and the other 3%. If he wants to earn $550 annual interest, how much should he invest at each rate?
Answer by SAT Math Tutor(36)

You need to set up 2 linear equations with variables x and y where x represents the amount invested in account 1 and y represents the amount represented in account 2.
x + y = 15000
x*.04 + y*.03 = 550
Solve the first equation for x:
x = 15000 - y
Plug this into the second equation:
(15000 - y)*.04 + y*.03 = 550
Then, solve:
600 - .04*y + y*.03 = 550
600 - .01*y = 550
-.01*y = -50
y = 5000
So, going back to what we have for x:
x = 15000 - 5000 = 10000
So y = 5000, x = 10000